what would the surface area of a rectangular prism with the dimensions of length: 6.5cm width: 15.1 cm height: 4.5 cm
Answer by awang1996(50)

what would the surface area of a rectangular prism with the dimensions of length: 6.5cm width: 15.1 cm height: 4.5 cm


surface area=the sum of the area of all faces
base+base+left+right+front+back
base=6.5*15.1=98.15
left=15.1*4.5=67.95
front=29.25
98.15+98.15+67.95+67.95+29.25+29.25=390.7

answer:390.7 cm cubed